If f(x)=x^2-3 and f(a)=46, what is the value of a? Explain.
Answer by lmeeks54(111)

The only tricky thing here is to understand what is meant by f(a) = ?
...
This is basically just solving the quadratic equation.
Given:
f(x) = x^2 - 3
f(a) = 46
...
In this case, f(a) means solving the f(x) equation as if a = x.
...
So, f(x) = f(a)
x^2 - 3 = 46
...
The easiest way to solve quadratic equations is to rearrange them (using the rules of algebra) so that the equation is set = 0
...
Therefore, subtract 46 from both sides and simplify:
x^2 - 3 - 46 = 46 - 46
x^2 - 49 = 0
...
Factor the quadratic equation:
x^2 - 49 = 0 is factored to become:
(x + 7)(x - 7) = 0
...
Solving for x to satisfy the factored quadratic equation:
x = 7, or -7
...
So, when f(a) = 46, a = 7 or -7
...
check the work:
(7)^2 - 3 = 46
49 - 3 = 46
46 = 46 checks
...
Note: if we substitute -7 into the function, we get the same answer.
a = 7, -7
...
cheers,
Lee